# This work was supported by project from DOE-CESER

### v1

@@ VII-A Lightning Arrestors
IEEE Std C62.11-2020 [190], IEEE Standard for Metal-Oxide Surge Arresters for AC Power Circuits (¿1 kV), states that before utilizing them in the field, the lightning arrestors are typically made to go through a process called thermomechanical conditioning by applying controlled thermal, mechanical, and electrical stress to them to improve their performance and reliability [190]. From Fig. 25, it can be observed that over a 24-hour period, these components would be undergoing thermal stress between $-40\degree C$ to $+60\degree C$. However, this testing does not account for temperatures lower than $-40\degree C$ and for longer durations. For example, an extreme cold event like the great Texas freeze in 2021 [190], led to very cold temperatures for up to 8 days. This standard also provides the ambient air temperature and arrestor temperature limits as shown in Tab. VI
@@ VII-F Relays
Typically, electromechanical direct current (DC) auxiliary relays, in hot conditions, pick up at higher proportions of the rated voltage (80% or less) compared to cold conditions (72% or less of rated voltage). For alternate current (AC) relays, they should be able to continuously withstand 110% of the rated voltage with maximum temperature rises as presented in Tab. 5 [151]. In Tab. 5 from [151], limits of temperature rise for coils for various insulation classes (105 C – 220 C) for varying ambient temperatures of both DC and alternate current (AC) type auxiliary relays are given.
@@ VIII DISCUSSION
In this section, we consolidate and discuss the information on different topics presented in this paper so far. These topics include the following: 1) A summary of the direct and indirect impact of extreme heat and cold on different equipment (categorized as primary and secondary), 2) A summary of different standards associated with different equipment and violation of equipment standards recommended operational temperature limits across U.S. geography, 3) Inter-dependency of equipment failures, and 4) Possible cascade failure due to equipment outage.
@@ VIII-A Summary of equipment impacts due to extreme heat and cold temperatures
Insulators and Bushings - Structural damage due to cycling of extreme temperatures Flashover Puncture due to flashover because of polymeric material
@@
Conductor Derating Thermal loading Ice and snow accumulation reduces right-of-way Galloping
@@
Lightening arrestor Manufacturer operational temperature - Manufacturer operational temperature -
@@ VIII-B Summary of standards and equipment operational temperature violations due to climate change
A summary of all the standards provided in this paper and potential modeling gaps present in the literature for each equipment has been presented in Tab. XI. For Example, in all the standards except for Circuit switchers and Circuit breakers, the permitted level of ice loading on the equipment has not been discussed, which would be critical information to know for the manufacturers or the equipment designers to account for ice loading due to extreme cold events such as the Polar vertex. Additionally, although the recommended maximum leakage rates for the instrument transformers at extremely cold conditions (up to $-50\degree C$) have been provided in [204], the corresponding recommended leakage rates at extremely high ambient temperatures (at higher than $40\degree C$) are not currently available in the literature.
@@
From Tab. XII, it can be observed that except for capacitor banks and the relays, the recommended ambient temperature limits for all the other equipment undergo violations in at least one of the considered two categories. Especially for battery and grounding resistor equipment, their recommended ambient temperature limits would be violated for all the considered extreme scenario anomaly conditions. This clearly shows a critical need to revisit and revise the current standards to ensure that the power system equipment performance is considered robust under all possible extreme weather conditions.
TABLE XII: Ambient temperature violations possibilities for various equipment based on existing standards
@@ VIII-C Inter-dependency of equipment
Beyond the direct effects of heat domes and polar vortexes on equipment, component failures can damage other parts of the power grid. This can result in cascading failures or damage that impacts grid reliability. In this section, some possible impacts of equipment on one another has been presented. However, these examples are not exhaustive and are only an indication of the interdependency of equipment in power grids.
Figs. 27(a), 27(b), and 28 demonstrate the impact of different equipment failures on two or three components each. These figures highlight each equipment and its corresponding failure mode due to extreme temperatures. The arrows indicate the impact of the equipment on the equipment of focus. For Example, in Fig. 27(a), the equipment of focus are the Substation and distribution transformer, which are highlighted by a black outline.
As shown in Fig. 27(a), many different equipment failures can affect some grid components. For Example, transformers are one of the most vulnerable components with inter-dependency; seven types of equipment failures can cause conditions that could lead to damage or failure of distribution transformers, and twelve types can do the same to substation transformers. Many of these dependencies occur in equipment that is either part of operation or protection, such as lightning arresters, grounding resistors, and relays that allow overcurrent or overvoltage to flow through the transformer. It is to be noted that the equipment considered to show interdependency is not exhaustive; it only serves as a demonstration of the interdependency of grid equipment and indication.
Similarly, Fig. 27(b) demonstrates the vulnerabilities of power lines (transmission and distribution conductors) and power line carrying equipment in transmission and distribution systems. Power lines can be impacted by numerous equipment throughout the grid, such as failures in circuit breakers, lightning arrestors, insulators, and capacitors. Most events that affect power lines cause or fail to protect the lines from overcurrent or overheating.
While Fig. 27(a) and Fig. 27(b) display the effects of some components on a few others, Fig. 28 shows a more interwoven chain of influence. Fig. 28 focuses on equipment with relationships to voltage regulators, frequency changers, and capacitor banks. Failures in components such as harmonic filters, volt-ampere reactive (VAR) devices, static VAR compensators (SVCs), and flexible alternating current transmission system (FACTS) devices cause the lifespans of capacitor banks to decrease. It is to be noted that not every grid infrastructure may have all the components mentioned above working together. However, it could be a combination of equipment for utility planning and operation requirements. Additionally, events in capacitor banks and voltage regulators can influence the distribution transformers.
@@
(a) Inter-dependency on transformers.
(b) Inter-dependency on power lines and power line carrier equipment.
Figure 27: Example of inter-dependency of different equipment in power grids.
@@ VIII-D Cascade Connections
While Figs. 27(a), 27(b), and 28 characterize many specific inter-dependency relationships, they do not capture the potential for cascading effects in power grids. Figs. 29(a) and 29(b) demonstrate a few possibilities for chains of equipment failures to occur under specific extreme weather conditions, although there are many possibilities that are not exhaustively presented in this work.
Fig. 29(a) identifies potential cascades in substations under extremely high temperatures. For Example, a heat-related failure in a relay can prevent a circuit breaker from tripping when a fault is present in the substation. This can cause overcurrent or overvoltage to flow through the substation, potentially damaging many other Substation components such as transformers, voltage regulators, and distribution busbars. These conditions could cause distribution busbars to overheat, damaging bus support insulators. Similar cascading effects are observable if the chain begins with the other weather-related equipment failure events in Fig. 29(a).
The contingency chain for transmission lines and carrier equipment in extreme cold and icing conditions is shown in Fig. 29(b). Cold ambient temperatures and icing can affect transmission equipment, leading to substation issues. For Example, towers that carry transmission lines can be brought down by the weight of ice accumulation. This will likely cause an open circuit, ground fault, or short circuit in the transmission conductors, which can lead to arcing in a circuit breaker or transformer. A transformer or circuit breaker failure could lead to power outages and damage to substation equipment.
(a) Example of cascades in Substation equipment during extreme heat.
(b) Example of cascades on transmission during extreme cold and icing.
Figure 29: Example of cascades in power grids during extreme weather conditions.
@@ IX-A Background Summary
In addition to the changes in the power grid described above, there is also a significant change in global climate that is impacting the overall average temperatures and extreme temperatures that the power grid needs to operate under. The past few decades have seen an increase in the frequency of occurrences and intensity of heat waves consistently indicating that the trend may not be changing in the near future. Not only the frequency, and intensity of the heat waves are increasing but also a more concerning aspect is that the duration of heat waves is also increasing. The overall hot conditions are leading to uncertain and relatively short duration of extreme cold scenarios owing to the phenomenon of disrupted polar vortex. This causes the cold winds to blow further south away from the poles bringing unexpected snow storms and severe ice conditions in normally hotter regions. The increased heat and increased cold in unexpected regions of the country have increased the vulnerability of the power grid equipment to heat domes, heat waves, polar vortexes, and severe ice conditions.
@@ IX-B Technical Conclusions
The present paper discusses the climate change phenomenon in a certain detail in the context of how the grid infrastructure is vulnerable to the changing climate. Therefore, this paper has discussed the most vulnerable set of T&D grid equipment including the various components in a substation providing the details of the individual equipment failure modes, its immediate impact on the other equipment, and is extended to cascaded failures of equipment. The failure modes and failure relation to other equipment are derived in the form of failure influence diagrams and cascaded failure influence diagrams. These diagrams give a better understanding of the overall vulnerability of the grid infrastructure. For each equipment, the paper also discusses existing and potential approaches that the grid planners and operators can take to monitor, measure, and mitigate the impacts of climate change (particularly, heat domes, polar vortexes, and icing). The paper also discusses the standards and their readiness in terms of the extreme temperatures experienced in the various regions of the different states.
@@ IX-C Future Work
While this work establishes the fundamental understanding of infrastructure impacts due to extreme temperatures from the literature review, industry articles, etc., it is important to translate this information into actionable steps regarding equipment availability, system contingency analysis, and overall grid stability analysis. We will expand on providing these insights along with gap analysis and research priorities in the future.

### v2

@@ VII-A Lightning Arrestors
IEEE Std C62.11-2020 [196], IEEE Standard for Metal-Oxide Surge Arresters for AC Power Circuits (¿1 kV), states that before utilizing them in the field, the lightning arrestors are typically made to go through a process called thermomechanical conditioning by applying controlled thermal, mechanical, and electrical stress to them to improve their performance and reliability [196]. From Fig. 25, it can be observed that over a 24-hour period, these components would be undergoing thermal stress between $-40\degree C$ to $+60\degree C$. For example, an extreme cold event like the great Texas freeze in 2021 [196], led to cold temperatures for prolonged period of time up to 8 days. Furthermore, this testing does not account for temperatures lower than $-40\degree C$ and also for longer durations. This standard also provides the ambient air temperature and arrestor temperature limits as shown in Tab. VI
@@ VII-F Relays
Typically, electromechanical direct current (DC) auxiliary relays, in hot conditions, pick up at higher proportions of the rated voltage (80% or less) compared to cold conditions (72% or less of rated voltage). For alternate current (AC) relays, they should be able to continuously withstand 110% of the rated voltage with maximum temperature rises as presented in Tab. 5 [149]. In Tab. 5 from [149], limits of temperature rise for coils for various insulation classes (class 105 – class 220) for varying ambient temperatures of both DC and alternate current (AC) type auxiliary relays are given.
@@ VIII DISCUSSION
Figure 27: Example of inter-dependency of different equipment in power grids: transformers.
Figure 28: Example of inter-dependency of different equipment in power grids: power lines and power line carrier equipment.
In this section, we consolidate and discuss the information on different topics presented in this paper so far. These topics include the following: 1) A summary of the direct and indirect impact of extreme heat and cold on different equipment (categorized as primary and secondary), 2) A summary of various equipment standards and violation of their recommended operational temperature limits across U.S. geography are presented to show the need for revision of these standards in light of heat waves and polar vortex, 3) Inter-dependency of equipment failures, and 4) Possible cascade failure due to equipment outage.
@@ VIII-A Summary of equipment impacts due to extreme heat and cold temperatures
Insulators and Bushings - Structural damage due to cycling of extreme temperatures Flashover due to icing and dust deposition Puncture due to flashover because of polymeric material
@@
Conductor Derating Thermal loading Ice and snow accumulation reduces right-of-way, Metal contraction which increases line tension causing damage to structure Galloping
@@
Lightning arrestor Manufacturer operational temperature - Manufacturer operational temperature -
@@ VIII-B Summary of standards and equipment operational temperature violations due to climate change
A summary of all the standards provided in this paper and potential modeling gaps present in the literature for each equipment has been presented in Tab. XI. For Example, in all the standards except for Circuit switchers and Circuit breakers, the permitted level of ice loading on the equipment has not been discussed, which would be critical information to know for the manufacturers or the equipment designers to account for ice loading due to extreme cold events such as the Polar vortex. Additionally, although the recommended maximum leakage rates for the instrument transformers at extremely cold conditions (up to $-50\degree C$) have been provided in [210], the corresponding recommended leakage rates at extremely high ambient temperatures (at higher than $40\degree C$) are not currently available in the literature.
@@
Equipment Standards Year Modeling Gaps
Lightning arrestors IEEE Std C62.11 2020 Thermal stress testing for longer than 24 hrs, Ice loading
Grounding resistors IEEE C57.32 2015 Cold spot temperature limits, resistance variance at extreme cold temperatures and Ice loading
Capacitor Banks IEEE Std 1036 2020 Ice loading
Harmonic filters IEEE Std 1531 2020 Ambient temperature limits at extreme cold conditions, Ice loading
Power fuses IEEE Std C37.41 2016 Ice loading
Relays IEEE Std C37.90, IEEE Std C37.91, IEEE Std C37.90.3 2005, 2021, 2023 Ice loading
Shunt reactors IEEE Std C57.21 2021 Ambient temperature limits at extreme cold conditions, Ice loading
Composite insulators IEEE Std 987 2001 Ice loading
Circuit switchers IEEE Std C37.30.1 2022 Current carrying capability temperature correction factor at extreme hot ambient temperatures
Circuit breakers IEEE Std C37.119, IEEE Std C37.04 2016, 2018 Adjustment needed for recommended ambient temperature limits for unusual service conditions
Overhead lines/ Underground cables IEEE Std 835, IEEE Std 835a, IEEE Std 738 1994, 2012, 2023 Recommended ambient temperature limits, Ice loading
Instrument transformers IEEE Std C57.13.5 2019 Leakage rates recommendations at extremely high ambient temperatures, Ice loading
Batteries IEEE Std 484, IEEE Std 485, IEEE Std 1635 2019, 2020, 2022 Recommendations for extreme hot/cold ambient temperatures, Ice loading
Distribution, Power and Auto Transformers IEEE Std C57.91, EEE Std C57.12.01, IEEE Std C57.12.00 2011,2020, 2021 Ice loading
@@
As mentioned earlier, the main reasoning behind considering cold conditions for a ”hot” state and vice-versa is because typically the manufacturers do not account for equipment’s operating conditions in hotter regions do not consider extreme cold conditions as they are an anomaly. In Tab,  XII,for example, for the considered five hottest regions, the historical extreme cold temperatures are binned into (-$20\degree C$, $30\degree C$) and ¡ $30\degree C$. We considered cold temperatures since the equipment’s operating conditions in hotter regions do not consider extreme cold conditions as they are an anomaly. However, the Texas great freeze is a good example of hot climatic region that experienced unexpected cold conditions. In the Tab,  XII, the standard for equipment such as power fuses says that their recommended standard ambient temperature ranges are between -$30\degree C$ and $40\degree C$ which aligns with observed cold temperatures between -$20\degree C$ and $30\degree C$ but does not align with extreme observed cold temperatures of ¡ -$30\degree C$. Additionally, from Tab. XII, it can be observed that except for capacitor banks and the relays, the ambient temperature limits from standards for all the other equipment undergo violations in at least one of the considered two anomaly categories in this work. Especially for battery and grounding resistor equipment, their recommended ambient temperature limits would be violated for all the considered extreme scenario anomaly conditions. This shows the need for revision of standards due to increased frequency, intensity, and duration of heat waves (Fig. 3) and polar vortex type events. This clearly shows a critical need to revisit and revise the current standards to ensure that the power system equipment performance is considered robust under all possible extreme weather conditions.
TABLE XII: Ambient temperature violation possibilities for various equipment based on existing standards. The values in the table represents if a particular equipment’s recommended ambient temperature range from standards is violated when compared with historical extreme cold/hot temperatures in hotter/colder regions in the U.S, respectively. Yes, No, and N/A in the table represents if there is a violation, no violation, and no available information respectively.
@@ VIII-C Inter-dependency of equipment
Beyond the direct effects of heat domes and polar vortices on equipment, component failures can damage other parts of the power grid. This can result in cascading failures or damage that impacts grid reliability. In this section, some possible impacts of equipment on one another has been presented. However, these examples are not exhaustive and are only an indication of the interdependency of equipment in power grids.
Figs. 27, 28, and 29 demonstrate the impact of different equipment failures on two or three components each. These figures highlight each equipment and its corresponding failure mode due to extreme temperatures. The arrows indicate the impact of the equipment on the equipment of focus. For Example, in Fig. 27, the equipment of focus are the Substation and distribution transformer, which are highlighted by a black outline.
As shown in Fig. 27, many different equipment failures can affect some grid components. For Example, transformers are one of the most vulnerable components with inter-dependency; seven types of equipment failures can cause conditions that could lead to damage or failure of distribution transformers, and twelve types can do the same to substation transformers. Many of these dependencies occur in equipment that is either part of operation or protection, such as lightning arresters, grounding resistors, and relays that allow overcurrent or overvoltage to flow through the transformer. It is to be noted that the equipment considered to show interdependency is not exhaustive; it only serves as a demonstration of the interdependency of grid equipment and indication.
Similarly, Fig. 28 demonstrates the vulnerabilities of power lines (transmission and distribution conductors) and power line carrying equipment in transmission and distribution systems. Power lines can be impacted by numerous equipment throughout the grid, such as failures in circuit breakers, lightning arrestors, insulators, and capacitors. Most events that affect power lines cause or fail to protect the lines from overcurrent or overheating.
While Fig. 27 and Fig. 28 display the effects of some components on a few others, Fig. 29 shows a more interwoven chain of influence. Fig. 29 focuses on equipment with relationships to voltage regulators, frequency changers, and capacitor banks. Failures in components such as harmonic filters, volt-ampere reactive (VAR) devices, static VAR compensators (SVCs), and flexible alternating current transmission system (FACTS) devices cause the lifespans of capacitor banks to decrease. It is to be noted that not every grid infrastructure may have all the components mentioned above working together. However, it could be a combination of equipment for utility planning and operation requirements. Additionally, events in capacitor banks and voltage regulators can influence the distribution transformers.
@@ VIII-D Cascade Connections
While Figs. 27, 28, and 29 characterize many specific inter-dependency relationships, they do not capture the potential for cascading effects in power grids. Figs. 30 and 31 demonstrate a few possibilities for chains of equipment failures to occur under specific extreme weather conditions, although there are many possibilities that are not exhaustively presented in this work.
Figure 30: Example of cascades in power grids during extreme weather conditions: substation equipment during extreme heat.
Figure 31: Example of cascades in power grids during extreme weather conditions: transmission during extreme cold and icing.
Fig. 30 identifies potential cascades in substations under extremely high temperatures. For Example, a heat-related failure in a relay can prevent a circuit breaker from tripping when a fault is present in the substation. This can cause overcurrent or overvoltage to flow through the substation, potentially damaging many other Substation components such as transformers, voltage regulators, and distribution busbars. These conditions could cause distribution busbars to overheat, damaging bus support insulators. Similar cascading effects are observable if the chain begins with the other weather-related equipment failure events in Fig. 30.
The contingency chain for transmission lines and carrier equipment in extreme cold and icing conditions is shown in Fig. 31. Cold ambient temperatures and icing can affect transmission equipment, leading to substation issues. For Example, towers that carry transmission lines can be brought down by the weight of ice accumulation. This will likely cause an open circuit, ground fault, or short circuit in the transmission conductors, which can lead to arcing in a circuit breaker or transformer. A transformer or circuit breaker failure could lead to power outages and damage to substation equipment.
@@ IX-A Background Summary
In addition to the changes in the power grid described above, there is also a significant change in global climate that is impacting the overall average temperatures and extreme temperatures that the power grid needs to operate under. The past few decades have seen an increase in the frequency of occurrences and intensity of heat waves consistently indicating that the trend may not be changing in the near future. Not only the frequency, and intensity of the heat waves are increasing but also a more concerning aspect is that the duration of heat waves is also increasing. The overall hot conditions are leading to uncertain and relatively short duration of extreme cold scenarios owing to the phenomenon of disrupted polar vortex. This causes the cold winds to blow further south away from the poles bringing unexpected snow storms and severe ice conditions in normally hotter regions. The increased heat and increased cold in unexpected regions of the country have increased the vulnerability of the power grid equipment to heat domes, heat waves, polar vortices, and severe ice conditions.
@@ IX-B Technical Conclusions
The present paper discusses the climate change phenomenon in a certain detail in the context of how the grid infrastructure is vulnerable to the changing climate. Therefore, this paper has discussed the most vulnerable set of T&D grid equipment including the various components in a substation providing the details of the individual equipment failure modes, its immediate impact on the other equipment, and is extended to cascaded failures of equipment. The failure modes and failure relation to other equipment are derived in the form of failure influence diagrams and cascaded failure influence diagrams. These diagrams give a better understanding of the overall vulnerability of the grid infrastructure. For each equipment, the paper also discusses existing and potential approaches that the grid planners and operators can take to monitor, measure, and mitigate the impacts of climate change (particularly, heat domes, polar vortices, and icing). The paper also discusses the standards and their readiness in terms of the extreme temperatures experienced in the various regions of the different states.
@@ IX-C Future Work
While this work establishes the fundamental understanding of infrastructure impacts due to extreme temperatures from the literature review, industry articles, etc., it is important to translate this information into actionable steps regarding equipment availability, system contingency analysis, and overall grid stability analysis. There are other important impacts of extreme cold and extreme heat related events on the power grid equipment like corrosion or structural integrity failures that lead to loss of equipment. There are other impacts that accelerate the end of life of equipment that may suddenly lead to acute failures eventually. We will expand on providing these insights along with gap analysis and research priorities in the future.